Parth (parthkohli):

How is \[ \lim_{x \to \infty} 1 - 5x^3 = -5\]?

OpenStudy (amistre64):

is x and z the same?

OpenStudy (anonymous):

ok divide by x^2

OpenStudy (anonymous):

after division by x^2 limit x->infinity 1/x^2-5 can you do it now?

OpenStudy (amistre64):

cubics dont converge

Parth (parthkohli):

Hmm, \(0 - 5\) if we do it individually.

Parth (parthkohli):

I got it!

OpenStudy (anonymous):

ok .

Parth (parthkohli):

But how are we allowed to divide by \(x^2\)?

OpenStudy (amistre64):

5x^3/x^2 = 5x tho ....

Parth (parthkohli):

And how did you get \(5x^3 \div x^2 = 5\)? It's 5x.

Parth (parthkohli):

Yeah

Parth (parthkohli):

It's the same if we divide by \(x^3\) anyway, just that the thing becomes 1/x^3

OpenStudy (anonymous):

does not matter you divided both if you take x^2 as LCM you get original back.

OpenStudy (anonymous):

no it wouldn't be same. always divide by highest power available in the question . here it is x^2

Parth (parthkohli):

Aha! Thank you! But I still don't get how we are allowed to divide both sides by x^3.

OpenStudy (anonymous):

sorry it is x^3

Parth (parthkohli):

\[ \lim_{x \to \infty} x^3({1 \over x^3} - 5)\] \[ x^3 \lim_{x \to \infty} -5 \]

Parth (parthkohli):

I think that you also factor the \(x^3\) out.

Parth (parthkohli):

Oh! Wait!

Parth (parthkohli):

\[ \lim_{x \to \infty}x^5 \times \lim_{x \to \infty}{1 \over x^3} - 5\]

Parth (parthkohli):

x^3*

Parth (parthkohli):

But isn't that the indeterminate form?

OpenStudy (amistre64):

why are you trying to makes sense of this? a cubic does not converge

Parth (parthkohli):

But it is given that the solution is -5.

OpenStudy (amistre64):

and i say the sky has floating cotton candy for clouds; does it make it true?

OpenStudy (anonymous):

no you cannot take it out always.just divide each term because if you take common then you will get x^3also then it will always gives you infinity.

Parth (parthkohli):

Well, may I know how you are allowed to divide by \(x^3\)? (I know that you divide the variable with the exponent as the highest power).

OpenStudy (apoorvk):

Uh oh. this should pretty simply be equal to negative infinity. Logically speaking.

Parth (parthkohli):

The original question is: \[ \lim_{z \to 0} {4z^2 + z^6 \over 1- z^3} \]

OpenStudy (amistre64):

if the original is that, why would you ask this other thing as a post?

Parth (parthkohli):

Infinity, not 0.

OpenStudy (anonymous):

@amistre64 you are right .it does not converges . this division rule works only when you have fractions .rational functions @ParthKohli you should take common x^3. for example

Parth (parthkohli):

Because they showed something like \[ \lim_{z \to \infty} {4z^2 + z^6 \over 1 - z^3} = {-\infty \over -5} \]

OpenStudy (anonymous):

the answer to your question is -infinity @ParthKohli

Parth (parthkohli):

Hmm, I was confused as to why it was -5 and not -infinity :P

Parth (parthkohli):

I mean... the question that I originally asked is supposed to be -infinity, right?

OpenStudy (anonymous):

yes right

OpenStudy (apoorvk):

Aye aye!